OpenStudy (anonymous):

Find the first three output values of the function f(z) = z2 – 1 + 2i. Use z = 0 as the first input value. –1 + 2i , –4 – 2i , 19 – 14i –1 + 2i , 2i , 3 + 2i –1 + 2i , –2 –+ 2i , –1 + 4i –1 + 2i , –4 – 2i , 11 + 18i

OpenStudy (perl):

f(0) = 0^2 -1 + 2i = -1 + 2i now use the output as the next input f(-1 + 2i) = (-1 + 2i ) ^2 -1 + 2i

OpenStudy (perl):

you can do this nicely with a TI 84/ TI 83 calculator

OpenStudy (perl):

console: 0 screen output: 0 type : ANS^2 -1 + 2i press enter , keep pressing enter

OpenStudy (anonymous):

i dont see the ANS button

OpenStudy (anonymous):

wait found it

OpenStudy (perl):

2nd (-)

OpenStudy (perl):

also make sure you can use i, which is 2nd . (next to 0 button)

OpenStudy (anonymous):

ok i clicked enter 4 times should i keep going? im still confused

OpenStudy (anonymous):

@perl ?

OpenStudy (perl):

you should see your sequence

OpenStudy (perl):

so first press 0. then type Ans^2 -1 + 2i Enter: -1 + 2i Enter : -4 -2i Enter : 11 + 18i

OpenStudy (perl):

if you keep pressing enter you see that it diverges , gets bigger and bigger

OpenStudy (perl):

Enter : -204 + 398i Enter: -116789 - 162382i etc

OpenStudy (perl):

this is a recursive sequence

OpenStudy (perl):

@Michele_Laino z0 = z0 z1 = f(z0) z2 = f(z1) = f(f(z0)) z3 = f(z2) = f(f(f(z0))) ...

OpenStudy (michele_laino):

ok! I understand!

OpenStudy (perl):

no problem :) the directions were not clear

OpenStudy (perl):

I believe this is called a 'Julia set'

OpenStudy (perl):

if I remember correctly given a complex number c, and a seed z0 f(z0)=z0 f(zn) = (z_n-1)^2 + c , iterating the function above produces a Julia set or Julia sequence or more simply f(z) = z^2 + c is the Julia set